# CPCL: Cross-Modal Prototypical Contrastive Learning for Weakly Supervised Text-based Person Re-Identification

### v1

@@ 4.3.4 Exploration of Momentum updating scheme for PMM
Figure 5: t-SNE visualization of 10 random person samples in CUHK-PEDES. Text and image samples are denoted as triangles and circles, respectively. Samples with the same color indicate that they belong to the same person. The same letter means that it corresponds to the same person.
@@
4.3.5 Two-stage OPLM vs One-stage OPLM
Although we propose a One-stage OPLM that can mine valuable unclustered samples, some image-text pairs still are unminable. Based on this fact, we propose a two-stage OPLM to make full use of these image-text pairs and conduct experiments to evaluate the effectiveness of the two-stage training scheme. As shown No.3 vs No.2 in Table 5, CPCL achieves better performance with the Two-stage OPLM training scheme. This phenomenon is intuitive, as the un-clustered image-text pairs are valuable and more training data allows the model to perform better on the test set.
4.4 Analysis of Qualitative Results
4.4.1 Visualization of Saliency Map
@@
To better understand visual and textual alignment, we use Grad-CAMSelvaraju et al. (2017) to compute the saliency map based on the last attention layer’s [CLS] token of the image encoder. The objective is the cosine similarity between the encoding of the input image and the text encoding. As shown in Fig. 4, our method has guided the model to learn the semantic relevance of visual the textual modalities. Compared with the baseline, CPCL could focus more on attributes of human bodies described by the text.
4.4.2 Visualization of t-SNE
We use t-SNEVan der Maaten and Hinton (2008) to visualize the embeddings of image and text samples belonging to 10 random persons in CUHK-PEDES. As shown in Fig. 5, our approach has better intra-class compactness and inter-class separability the the baseline. The compactness and separability contribute to better weakly supervised TPRe-ID performance. We also can find that text samples are more diverse compared to image samples. This phenomenon is consistent with the reality of the data, as annotators describe the same image differently. This finding may indicate that the difficulty of the weakly supervised TPRe-ID lies in how to further process the text samples.
4.4.3 Visualization of retrieval results
@@
We showcase the top-10 retrieval results obtained using textual descriptions from both the baseline and CPCL. As shown in Fig. 6, CPCL achieves more accurate retrieval results for a textual query. However, it is worth noting that our model also returned an incorrect retrieval result. This occurrence may be attributed to the global matching method employed in our model, which directly aligns global visual and textual embeddings and might overlook some crucial details. In future work, we are planning to alleviate this issue.
5 Conclusion
CPCL has shown considerable improvements on weakly supervised TPRe-ID. The core of the CPCL is to bring instance-level samples closer to their prototypes. Though the results are positive, there still exists a large gap from the fully supervised TPRe-ID, suggesting that the label by the clustering algorithm is not satisfactory enough. We argue that conducting pseudo labels refinery can further improve the performance of the model and we will further investigate it in our future work. Moreover, we hope that our proposed CPCL could inspire and facilitate follow-up works with potential.

### v2

@@ Appendix B The Number of Outliers during Training
Appendix C Analysis of Qualitative Results
C.1 Visualization of t-SNE
We employ t-SNE(Van der Maaten and Hinton 2008) to visualize the embeddings of image and text samples from 10 randomly selected individuals in the CUHK-PEDES dataset. As depicted in Fig. 6, our approach demonstrates superior intra-class compactness and inter-class separability compared to the baseline. This enhanced compactness and separability are instrumental in achieving better weakly supervised text-based person retrieval performance.
The improved intra-class compactness ensures that samples from the same person are clustered more tightly, while the enhanced inter-class separability ensures that samples from different individuals are more distinctly separated. This dual improvement not only facilitates the model’s ability to distinguish between different individuals but also enhances its robustness in recognizing individuals under weak supervision.
Additionally, our analysis reveals that text samples exhibit greater diversity compared to image samples. This observation aligns with the inherent nature of the data, where annotators often describe the same image in various ways. This diversity in textual descriptions suggests that the challenge of weakly supervised text-based person retrieval may lie in effectively processing and leveraging these varied text samples.
C.2 Visualization of Saliency Map
@@
To gain deeper insights into the alignment between visual and textual modalities, we employ Grad-CAM(Selvaraju et al. 2017) to compute the saliency map based on the [CLS] token of the last attention layer in the image encoder. The objective of this analysis is to measure the cosine similarity between the encoding of the input image and the text encoding.
As illustrated in Fig. 7, our method effectively guides the model to learn the semantic relevance between visual and textual modalities. This is evident in the model’s ability to focus more on the attributes of human bodies described by the text, as compared to the baseline. This enhanced focus on textual attributes suggests that our approach not only aligns the visual and textual data but also enriches the model’s understanding of the semantic context provided by the text.
Furthermore, this improved alignment can be attributed to the model’s ability to capture nuanced relationships between visual features and textual descriptions. By leveraging the saliency maps, we can observe how the model prioritizes certain visual elements based on the textual input, which is crucial for tasks that require a high degree of semantic understanding, such as image captioning or visual question answering.
C.3 Visualization of retrieval results
We present the top-10 retrieval results obtained using textual descriptions from both the baseline and CPCL (Cross-Prototype Classification Learning). As illustrated in Fig. 8, CPCL demonstrates superior accuracy in retrieving relevant images for a given textual query. This improvement is attributed to the model’s enhanced ability to align visual and textual modalities more effectively.
@@ C.3 Visualization of retrieval results
However, it is important to note that our model also returned an incorrect retrieval result. This occurrence can be attributed to the global matching method employed in our model, which directly aligns global visual and textual embeddings. While this approach offers computational efficiency, it may inadvertently overlook some crucial local details that are essential for accurate retrieval.
This observation highlights a potential limitation of our current model and suggests that future work should focus on refining the matching strategy. One promising direction is to incorporate local feature matching, which could provide a more nuanced understanding of the visual and textual content. By combining global and local information, we aim to enhance the model’s ability to capture fine-grained details, thereby improving overall retrieval accuracy.
No. Methods CUHK $\rightarrow$ ICFG
R@1 R@5 R@10
Fully Supervised Text-based Person Retrieval
1 MIA (Niu et al. 2020) 19.35 36.78 46.42
2 SCAN (Lee et al. 2018) 21.27 39.26 48.83
3 SSAN (Ding et al. 2021) 29.24 49.00 58.53
4 LGUR (Shao et al. 2022) 34.25 52.58 60.85
5 IRRA(Jiang and Ye 2023) 42.40 62.10 69.61
6 RaSa(Bai et al. 2023) 50.59 67.46 74.09
Weakly Supervised Text-based Person Retrieval
7 CMMT(Zhao et al. 2021) 31.16 50.02 17.31
8 CPCL(Ours) 39.66 59.01 67.67
Table 8: Comparison with other methods on domain generalization task. We adopt CUHK-PEDES and ICFG-PEDES as the source domain and the target domain in turn.
No. Methods ICFG $\rightarrow$ CUHK
R@1 R@5 R@10
Fully Supervised Text-based Person Retrieval
0 Dual Path (Zheng et al. 2020) 7.63 17.14 23.52
1 MIA (Niu et al. 2020) 10.93 23.77 32.39
2 SCAN (Lee et al. 2018) 13.63 28.61 37.05
3 SSAN (Ding et al. 2021) 21.07 38.94 48.54
4 LGUR (Shao et al. 2022) 25.44 44.48 54.39
5 IRRA(Jiang and Ye 2023) 33.45 56.29 66.31
6 RaSa(Bai et al. 2023) 50.70 72.40 79.58
Weakly Supervised Text-based Person Retrieval
7 CMMT(Zhao et al. 2021) 33.07 54.24 63.99
8 CPCL(Ours) 35.48 57.52 66.57
Table 9: Comparison with other methods on domain generalization task. We adopt ICFG-PEDES and CUHK-PEDES as the source domain and the target domain in turn.
Appendix D Domain Generalization
We perform a series of domain generalization experiments to assess the generalization ability of CPCL. Specifically, we evaluate the model’s performance on a target domain using a model trained on a source domain, as detailed in Table 8 and Table 9.
These experiments are designed to test how well CPCL can adapt to new, unseen domains without additional training. By transferring the model from a source domain to a target domain, we aim to understand the extent to which CPCL can generalize its learned representations across different datasets. The experiments demonstrate that our model achieves commendable performance. While it does not yet match the performance of state-of-the-art supervised methods, we are pleasantly surprised to find that our approach outperforms many existing supervised methods. The fact that our weakly supervised model can surpass several supervised methods suggests that the underlying principles and techniques we have developed are highly effective.
This finding also raises intriguing questions about the potential of weakly supervised learning in domains traditionally dominated by fully supervised approaches. There are three possible answers to this phenomenon: one is that supervised methods suffer from overfitting problems, two is that the labels in the training data are noisy, and three is that both exist. Since this problem is beyond the scope of this paper, we intend to explore the boundaries of this problem in subsequent work.
Our domain generalization experiments not only validate the robustness of CPCL but also open up new research directions for enhancing cross-domain adaptability, which is essential for deploying machine learning models in real-world scenarios.
Appendix E Impact of Clustering Hyper-parameters
We use DBSCAN(Ester et al. 1996) algorithm and Jaccard distance(Zhong et al. 2017) with $k$-reciprocal nearest neighbors for clustering images and texts separately, where the $k=20$. As for images, the maximum distance $d_{v}$ between neighbors is set as 0.5 and the minimal number of neighbors $neb_{v}$ for a dense point is set as 2. As for texts, the maximum distance $d_{t}$ between neighbors is set as 0.6 and the minimal number of neighbors $neb_{t}$ for a dense point is set as 4.
The maximum distance and minimal number of neighbors for a dense point are hyper-parameters of the DBSCAN algorithm, which affect the final number of clusters. In general, if the $d_{v}$ and $d_{t}$ is chosen too large, the clusters will merge and it is possible that a cluster will contain a very large number of samples. Conversely, if they are set very small, then most of the samples will be considered as outliers. As for the minimal number of neighbors $neb_{v}$ and $neb_{t}$, the larger they are, the more samples will be considered as outliers. Conversely, if the value of they is too small, then there will be a number of core points, which in turn will lead to multiple clusters being merged into one.
We analyze the influence of them on the CUHK-PEDES as shown in Table 10 and Table 11. We find that the image samples are more sensitive to different hyper-parameters, as No.2 and No.5 shown in Table 10. On the contrary, the text samples are more robust to different hyper-parameters, and the performance of the OPLM fluctuates very little for different parameters. We conjecture that this is related to the number of samples in the different modalities. As for the CUHK-PEDES dataset, each image will typically correspond to two text samples, so the number of samples for text is approximately twice as large as for images.
No. $d_{v}$ $neb_{v}$ CUHK-PEDES
R@1 R@5 R@10
0 0.3 2 70.31 86.84 92.00
1 0.5 70.03 87.28 91.78
2 0.7 60.58 79.18 86.08
3 0.5 2 70.03 87.28 91.78
4 4 69.51 87.09 91.76
5 8 65.03 83.92 89.98
Table 10: Performances of CPCL with different of $d_{v}$ and $neb_{v}$, where $d_{v}$ represents the maximum distance between two image samples and $neb_{v}$ represents the minimal number of image neighbors for an image dense point in DBSCAN. $d_{t}=0.6$ and $neb_{t}=4$.
No. $d_{t}$ $neb_{t}$ CUHK-PEDES
R@1 R@5 R@10
0 0.4 4 68.97 86.06 91.16
1 0.6 70.03 87.28 91.78
2 0.8 69.75 86.34 91.47
3 0.6 2 70.04 86.40 91.99
4 4 70.03 87.28 91.78
5 8 70.42 86.55 91.78
Table 11: Performances of CPCL with different of $d_{t}$ and $neb_{t}$, where $d_{t}$ represents the maximum distance between two text samples and $neb_{t}$ represents the minimal number of text neighbors for a text dense point in DBSCAN. $d_{v}=0.5$ and $neb_{v}=2$.
Appendix F Implementation Details
The Hyper-parameters in training
The batch size is set to 128, and the model is trained with Adam optimizer(Kingma and Ba 2015) for 60 epochs with a learning rate initialized to $1\times 10^{-5}$ and cosine learning rate decay. In the beginning, the 5 warm-up epochs linearly increase the learning rate from $1\times 10^{-6}$ to $1\times 10^{-5}$. Following the clustering-based method CMMT(Zhao et al. 2021), we use DBSCAN(Ester et al. 1996) for clustering before each epoch. All experiments are performed on a single RTX3090 24GB GPU.
The Text Encoder of CLIP. In this study, we adopt a CLIP text encoder to extract the text representation, which is a Transformer(Vaswani et al. 2017) model that has been modified by Radford et al.(Radford et al. 2021). Following the CLIP, we first employ lower-cased byte pair encoding (BPE) with a vocabulary size of 49152(Sennrich et al. 2015) to tokenize the input text description. To demarcate the beginning and end of the sequence, we enclose the tokenized text $\{f^{t}_{sos},f^{t}_{1},\ldots,f^{t}_{eos}\}$ with special [SOS] and [EOS] tokens. Subsequently, these tokenized text representations are fed into the transformer and capture correlations among each path by masked self-attention. Finally, we obtain the global text representation by linearly projecting the last layer of the transformer at the [EOS] token $f^{t}_{eos}$ into the image-text joint embedding space.
The Image Encoder of CLIP. CLIP image encoder is a ViT(Dosovitskiy et al. 2020) model. Given an input image $I\in R^{H\times W\times C}$, where H, W, and C denote its height, width, and number of channels, respectively. In order to input Image $I$, we first split $I$ into $N=H\times W/P^{2}$ fixed-sized patches $\{f^{v}_{i}|i=1,2,\textperiodcentered\textperiodcentered\textperiodcentered,N\}\in R^{1\times D}$, where N, P and D denote number, size and dim of the patch after a trainable linear projection. Spatial information is incorporated by adding learnable position embeddings and an extra learnable [CLS] embedding token denoted as $I_{cls}$ is prepended to the input patch token sequences $\{f^{v}_{cls},f^{v}_{1},\ldots,f^{v}_{N}\}$. Then the patch token sequences are input into L-layer transformer blocks to model correlations of each patch token. Finally, a linear projection is adopted to map $f^{v}_{cls}$ to the joint image-text embedding space, which serves as a global image representation.
Appendix G Details of Dataset
G.1 CUHK-PEDES
The CUHK-PEDES(Li et al. 2017) dataset represents a pioneering effort in text-to-image person retrieval, encompassing 40,206 images and 80,412 textual descriptions associated with 13,003 individual identities. The training set comprises 11,003 identities, 34,054 images, and 68,108 textual descriptions, while the validation and test set each includes 3,078 images and 6,158 textual descriptions, respectively, representing a total of 1,000 unique identities in each set.
G.2 ICFG-PEDES
ICFG-PEDES(Ding et al. 2021) is a dataset proposed in 2021, built upon the MSMT17(Wei et al. 2018) dataset. It is unique in that each pedestrian image is associated with only one text description. The dataset consists of 54,522 pedestrian images and an equal number of fine-grained text descriptions, covering a total of 4,102 individual pedestrians. As per official settings, the data is divided into a training set and a test set. The training set comprises 34,674 image-text pairs, corresponding to 3,102 pedestrian IDs, while the test set contains 19,848 image-text pairs, involving 1,000 pedestrian IDs. While this dataset has a smaller number of text descriptions compared to the CUHK-PEDES dataset, the descriptions provided are more detailed and fine-grained.
G.3 RSTPReid
The RSTPReid(Zhu et al. 2021) dataset, comprises 20,505 images and 41,010 corresponding text descriptions, relating to a total of 4,101 unique pedestrians. The images were captured from 15 cameras, enhancing the dataset’s realism by replicating real-world scenarios. Notably, this dataset guarantees that each image is associated with precisely two text descriptions, and each pedestrian is depicted by five images captured from multiple cameras. The dataset is thoughtfully divided into training, verification, and test sets, as per official guidelines. The training set comprises 18,505 images, 37,010 text descriptions and 3,701 individual pedestrians. Both the test and training sets include 1,000 images and 2,000 descriptions, collectively representing 200 distinct pedestrians.